OpenStudy (anonymous):

A,B and C are three collinear points.A and B are respectively (3,4) and (7,7) and AC is equal to 10 units?

OpenStudy (anonymous):

a sry, i missunderstud you. I made bc = 10, :)

OpenStudy (anonymous):

C(x,y) = (3,4) + (k4,k3) sqrt(16k^2+9k^2) = 10 sqrt(25k^2 )= 5k =10 k=2 so:C(x,y) = (3,4) + (2*4,2*3) = (3+8,4+6) = (11,10)

OpenStudy (anonymous):

@aftab13